Two years ago ,father was thee times as his son and two years hence,twice his age will be equal to five times that of his son.find their present ages
Answer by ankor@dixie-net.com(22740)

Let f = father's present age
let s = son's age
:
Write an equation for each statement, simplify as much as possible:
;
Two years ago, father was thee times as his son and
f - 2 = 3(s-2)
f - 2 = 3s - 6
f = 3s - 6 + 2
f = 3s - 4
:
two years hence, twice his age, will be equal to five times that of his son.
2(f + 2) = 5(s + 2)
2f + 4 = 5s + 10
2f = 5s + 10 - 4
2f = 5s + 6
Replace f with (3s-4) from the 1st statement equation
2(3s - 4) = 5s + 6
6s - 8 = 5s + 6
6s - 5s = 6 + 8
s = 14 yrs is the son's age
then
f = 3(14) - 4
f = 38 yrs is the father's age
:
You can see if that checks out in both statements